Sales $4,100.00
Operating costs excluding depreciation 3,053.00
EBITDA $1,047.00
Depreciation 300.00
EBIT $747.00
Interest 170.00
EBT $577.00
Taxes (40%) 230.80
Net income $346.20

Looking ahead to the following year, the company's CFO has assembled this information:

  • Year-end sales are expected to be 4% higher than $4.1 billion in sales generated last year.
  • Year-end operating costs, excluding depreciation, will equal 80% of sales.
  • Depreciation costs are expected to increase at the same rate as sales.
  • Interest costs are expected to remain unchanged.
  • The tax rate is expected to remain at 40%.

On the basis of this information, what will be the forecast for Edwin's year-end net income? Enter your answers in millions. For example, an answer of $10,550,000 should be entered as 10.55. Enter all values as positive numbers. Do not round intermediate calculations. Round your answers to two decimal places.

(in millions of dollars)
Sales $  
Operating costs excluding depreciation    
EBITDA $  
Depreciation    
EBIT $  
Interest    
EBT $  
Taxes    
Net income $  
Answer #1

Preaparing the Income Statement:-

Particular Calculations Amount in millions of $
Net Sales =4100*(1+4%)                         4,264.00
Less: Costs of Goods Sold =4264*80%                         3,411.20
EBITDA                            852.80
Less: Depreciation =300*(1+4%)                            312.00
EBIT                            540.80
Interest Expenses Unchanged                            170.00
Earning before tax                            370.80
Taxes @40%                            148.32
Net Income                            222.48

So, Forecasted year end Net Income = $222.48 millions